(D2 -2D +3)y = x2 -1


1
Expert's answer
2021-07-13T05:15:50-0400
"(D^2-2D+3)y=x^2-1"

The homogeneous differential equation


"(D^2-2D+3)y=0"

The characteristic equation


"r^2-2r+3=0"

"r_1=1-\sqrt{2}i,r_2=1+\sqrt{2}i"

The general solution of the homogenepus equation is


"y_h=c_1e^x\cos(\sqrt{2}x)+c_2e^x\sin(\sqrt{2}x)"

Find the particular solution of the nonhomogeneous differential equation


"y_p=Ax^2+Bx+C"

"y_p '=2Ax+B"


"y_p''=2A"

Substitute


"2A-4Ax-2B+3Ax^2+3Bx+3C=x^2-1"

"3A=1=>A=\dfrac{1}{3}"

"-4A+3B=0=>B=\dfrac{4}{9}"

"2A-2B+3C=-1=>C=-\dfrac{7}{27}"

Then


"y_p=\dfrac{1}{3}x^2+\dfrac{4}{9}x-\dfrac{7}{27}"

The general solution of the given nonhomogenepus equation is


"y=c_1e^x\cos(\sqrt{2}x)+c_2e^x\sin(\sqrt{2}x)+\dfrac{1}{3}x^2+\dfrac{4}{9}x-\dfrac{7}{27}"